Molly is making a steak recipe that calls for 4/5 cups of salt.
A) How much would be needed to make 1/3 of the recipe?
B) 3 and 1/2 of the recipe?
Show work.

Answer by jorel555(1290)

A) 4/5 x 1/3=4/15 cups of salt.
B) 4/5 x 7/2=2-4/5 cups of salt. ☺☺☺☺